let Tr be the rth term of an AP whose first term is a abd common difference is d is some positive integers m ,n ,m not equal to n , Tm=1÷n , Tn =1÷m then find the value of a-d

Grade:12th pass

1 Answers

Arun
5 years ago
Tm = 1/n
Hence
a + (m-1) d = 1/n
Similarly
a + (n-1) d = 1/m
Now solve both the equations
d = 1/mn
Hence a = 1/mn
Now a-d = 0